OpenStudy (anonymous):

Find the equation of the circle with center (1,3) and radius of 2. Please help me!

OpenStudy (anonymous):

@GretaKnows

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

this is the general form of a circle with center (a,b) and radius r (x - a)^2 + (y - b)^2 = r^2 now just plug in your values and you have the answer

OpenStudy (anonymous):

so would it look like this? (x-1)^2 + (y-3)^2 = 2^2? @welshfella

OpenStudy (welshfella):

yes